OpenStudy (mallorysipp234):

The sum of two numbers is 55 and the difference is 25 . What are the numbers?

hartnn (hartnn):

so let the 2 numbers be x and y. "sum of 2 numbers is 55" how will you make an equation in x and y using this info?

OpenStudy (mallorysipp234):

well x+y=55, difference is obviously subtraction so it would be something like x+(y-25)=55?

hartnn (hartnn):

the difference between x and y is just x-y so x-y = 25 makes sense?

OpenStudy (mallorysipp234):

right, but it says the sum of the two.

hartnn (hartnn):

`sum of two numbers is 55` -> x+y = 55 `he difference is 25` -> x-y = 25 you still have confusion?

OpenStudy (mallorysipp234):

ok i see

hartnn (hartnn):

now add both the equations, you will eliminate y and get x

OpenStudy (mallorysipp234):

40 and 15

hartnn (hartnn):

good! :)